# appendicesapc Appendix Table of Contents

### v1

@@ Appendix C DRPO Algorithm
Appendix D Construction the Ranking Preference Data
HH Dataset. To construct ranking preference data based on HH dataset, we first supervised fine-tune(SFT) the Qwen1.5-4B model with HH dataset. Subsequently, we sample $K=6$ responses for each prompt $x^{i}$ in HH dataset using trained SFT model (Qwen1.5-4B) with parameter $temperature=0.7,top\_k=40$. We combine the sampled data and the original pairwise data to obtain the $K=8$ responses. We then score each response with a reward model $R$. In our work, we use the RM-Deberta-v3-large-v2footnotemark: model as our reward model, which is based on the DeBERTaV3 [He et al., 2021], subsequently trained on the HH dataset.
Specifically, following Liu et al. [2024], we compute the score $s^{i}_{j}$ of each response $y^{i}_{j}$ with respect to the prompt $x^{i}$ as follows:
$s^{i}_{j}=\frac{1}{K}\sum_{l=1}^{K}\frac{e^{R(x^{i},y^{i}_{j})}}{e^{R(x^{i},y^{i}_{j})}+e^{R(x^{i},y^{i}_{l})}},\text{where}~{}R~{}\text{is a reward model}.$
Based on the discussion above, we can construct our ranking preference dataset as $D=\{x^{i},\mathbf{y}^{i},\mathbf{s}^{i}\}_{i=1}^{N}$, where $\mathbf{y}^{i}=(y^{i}_{1},...,y^{i}_{K})$ and $\mathbf{s}^{i}=(s^{i}_{1},...,s^{i}_{K})$. To support conducting an ablation study on the size $K$ of the ranking preference dataset, we also split this dataset of size $K=8$ into three subsets, each with sizes $K=2$, $K=4$, and $K=6$, respectively. We also provide an example of our ranking preference dataset as shown in Table 8.
UltraFeedback Dataset. This dataset itself contains four responses and corresponding scores, which can be directly used to construct a ranking preference dataset of size $K=4$. We simply normalize these scores to the range $[0,1]$.
VLFeedback Dataset. Similar to the UltraFeedback Dataset, this dataset can be used directly as a ranking preference dataset of size $K=4$. We simply normalize these scores to the range $[0,1]$.
Appendix E Experimental Details
Software and Hardware. We conduct our experiments on servers equipped with NVIDIA A6000 GPUs (48GB VRAM) and NVIDIA L20 GPUs (48GB VRAM), with NVIDIA CUDA Toolkit version 11.8. All experiments are implemented in Python 3.10.13 using the PyTorch 2.1.2 framework.
Training Qwen1.5 on HH Dataset. For the training of all qwen1.5 series models on the HH dataset, we implement our methods and baselines based on the alignment-handbook repository [^4] https://github.com/huggingface/alignment-handbook and design our Trainer based on trl repository [^5] https://github.com/huggingface/trl. In our experiments, all models are trained for one epoch using the Rmsprop optimizer [Ruder, 2016]. We linearly warm up the learning rate from 0 to $5\cdot 10^{-7}$ over 150 steps and use a batch size of 4. We performed the training on the entire training split, setting the maximum prompt length to 512 and the maximum length to 1024.
Training Mistral on UltraFeedback Dataset. To reduce memory consumption, we utilize Qlora[Dettmers et al., 2023] to facilitate the training of the Mistral model UltraFeedback dataset. We directly use the Mistral-sft-beta model [^6] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta that was trained with supervised fine-tuning on the UltraFeedback dataset as our SFT model. Similarly, we also use the Rmsprop [Ruder, 2016] optimizer in conjunction with DeepSpeed ZeRO stage 2 [Rasley et al., 2020] for one epoch, and apply a cosine learning rate schedule with a warm up ratio of 0.1. we use a global batch size of 32 with a learning rate $5\cdot 10^{-6}$. We further evaluate our method against recent preference alignment techniques, including ORPO and R-DPO [Chowdhury et al., 2024, Hong et al., 2024]. For these baselines, we utilize their publicly reported test results.
Training Qwen-VL-Chat on VLFeedback Dataset. For the training of Qwen-VL-Chat model on the VLFeedback dataset, similar to previous settings, we use the Rmsprop optimizer and a global batch size of 32. To facilitate efficient training and reduce memory consumption, we utilize LoRA tuning. We set the maximum prompt length to 640 and the maximum response length to 1280, as the response lengths in VLFeedback are longer than in other datasets. We use a cosine learning rate schedule with a warm-up ratio of 0.1 and a peak learning rate of $5\times 10^{-6}$.
Loss Function Parameters. Unless otherwise specified, we configure our diffNDCG loss function with default settings of $\gamma=0.9999$, and $\beta=1$. For the methods DPO, DPOBT, and LiPO, the parameter $\beta$ is consistently set at $0.1$ in HH dataset experiments.
Table 8: An example of our constructed ranking preference dataset based on HH dataset.
Prompt
@@ Appendix E Experimental Details
Response 1
You can learn to do magic card tricks by watching instructional videos or reading books about card tricks. You can also try practicing by performing tricks on your friends and family. If you want to learn more advanced tricks, you can take a course in magic.
Score: 0.84
Response 2
You can learn magic card tricks with a deck of cards and a little bit of practice. Here’s a simple trick:Deal 2 cards face down on the table, and 2 cards face up.Look at the 2 cards face up, and tell the other person which card you think is the “trick card”. Then, turn over the 2 remaining face down cards. If the 2 cards match the trick card, then you did the trick!
Score: 0.65
Response 3
You can learn magic tricks by reading books on magic and practicing the tricks you learn.
Score: 0.61
Response 4
Wow! Well, you can find magic tricks pretty easily online. How about this website here? It has some popular tutorials, and I can find some others if you like. Also, I can just share some general tips about how to improve at magic tricks: the most important thing to learn is how to “palm” or conceal your moves from your audience. What this means is, if you’re doing some kind of card trick or sleight-of-hand move where you hold a card and pretend to do something but actually drop the card while you’re grabbing for another one, your audience might notice that something is amiss if you hold the card in front of your hands the entire time. So if you can learn how to keep your hands and other objects in front of your face, it will make your tricks look more natural. There are a bunch of online guides about this, some of them here, and others you can find through your favorite search.
Score: 0.56
Response 5
Magic card tricks are a little different from card games like poker or bridge. Instead of betting on the outcome of a hand, card tricks involve the manipulation of the cards themselves. You can actually do card tricks with ordinary playing cards, but it’s best to use specially-made “magic cards” that have a special “magic” effect.
Score: 0.47
Response 6
Can you tell me what kind of magic you’re interested in? Are you interested in card tricks?
Score: 0.39
Response 7
Here are some basic magic card tricks that are easy to learn.
Score: 0.26
Response 8
That’s a tough question! Do you want to do them with friends, or are you looking to learn some useful techniques for interacting with people you know?
Score: 0.19
Prompt
@@
Response 1
You can learn to do magic card tricks by watching instructional videos or reading books about card tricks. You can also try practicing by performing tricks on your friends and family. If you want to learn more advanced tricks, you can take a course in magic.
Score: 0.84
Response 2
You can learn magic card tricks with a deck of cards and a little bit of practice. Here’s a simple trick:Deal 2 cards face down on the table, and 2 cards face up.Look at the 2 cards face up, and tell the other person which card you think is the “trick card”. Then, turn over the 2 remaining face down cards. If the 2 cards match the trick card, then you did the trick!
Score: 0.65
Response 3
You can learn magic tricks by reading books on magic and practicing the tricks you learn.
Score: 0.61
Response 4
Wow! Well, you can find magic tricks pretty easily online. How about this website here? It has some popular tutorials, and I can find some others if you like. Also, I can just share some general tips about how to improve at magic tricks: the most important thing to learn is how to “palm” or conceal your moves from your audience. What this means is, if you’re doing some kind of card trick or sleight-of-hand move where you hold a card and pretend to do something but actually drop the card while you’re grabbing for another one, your audience might notice that something is amiss if you hold the card in front of your hands the entire time. So if you can learn how to keep your hands and other objects in front of your face, it will make your tricks look more natural. There are a bunch of online guides about this, some of them here, and others you can find through your favorite search.
Score: 0.56
Response 5
Magic card tricks are a little different from card games like poker or bridge. Instead of betting on the outcome of a hand, card tricks involve the manipulation of the cards themselves. You can actually do card tricks with ordinary playing cards, but it’s best to use specially-made “magic cards” that have a special “magic” effect.
Score: 0.47
Response 6
Can you tell me what kind of magic you’re interested in? Are you interested in card tricks?
Score: 0.39
Response 7
Here are some basic magic card tricks that are easy to learn.
Score: 0.26
Response 8
That’s a tough question! Do you want to do them with friends, or are you looking to learn some useful techniques for interacting with people you know?
Score: 0.19
Appendix F GPT-4 Evaluation Details
We use gpt-4-1106-preview for all our experiments. Table 9 presents our template for requesting GPT-4 during our GPT-4 evaluation.
Table 9: The GPT-4 evaluation template. Text highlighted in yellow-orange represents the prompt; text in green represents the responses under comparison; text in blue represents the GPT-4 response, respectively.
User prompt
Here is a conversation between Human and Assistant.
Human:How can I get better at disagreeing with someone without causing an argument?Assistant:This is a difficult skill to learn. But there are ways you can do this, and I will help you.Human:How can you help me?Assistant:You can train in front of a mirror, and follow these steps. What I recommend is that you role-play with yourself, to practice and get comfortable with this skill. But please do this in a safe space. It’s important to practice safety. It’s also important to tell me if you would like help or instructions.Human:How can I practice this skill alone?’
Assistant: can now choose from different responses.
RESPONSE 1: You can also practice this in a safe environment. You can do this with a friend, a partner, or a therapist, and have a fun time practicing and getting comfortable with this skill. You can also do this with family, as this can be a great bonding experience. It’s also a good idea to practice in a social setting, like a group setting, to get comfortable with this skill.
END OF RESPONSE 1
@@ Appendix F GPT-4 Evaluation Details
END OF RESPONSE 2
Between these two responses, which response is more helpful, harmless, and concise? Pick one.
GPT-4 Response
Response 2 is more helpful, harmless, and concise. It directly addresses the question about practicing the skill alone by suggesting a specific role-play scenario. It provides a clear example of how to disagree without causing an argument, which is precisely what the human asked for. It also avoids suggesting potentially more complex or less safe scenarios involving other people, as in Response 1, which could lead to real disagreements or conflicts if not handled carefully.
User prompt
Here is a conversation between Human and Assistant.
Human:How can I get better at disagreeing with someone without causing an argument?Assistant:This is a difficult skill to learn. But there are ways you can do this, and I will help you.Human:How can you help me?Assistant:You can train in front of a mirror, and follow these steps. What I recommend is that you role-play with yourself, to practice and get comfortable with this skill. But please do this in a safe space. It’s important to practice safety. It’s also important to tell me if you would like help or instructions.Human:How can I practice this skill alone?’
Assistant: can now choose from different responses.
RESPONSE 1: You can also practice this in a safe environment. You can do this with a friend, a partner, or a therapist, and have a fun time practicing and getting comfortable with this skill. You can also do this with family, as this can be a great bonding experience. It’s also a good idea to practice in a social setting, like a group setting, to get comfortable with this skill.
END OF RESPONSE 1
@@
END OF RESPONSE 2
Between these two responses, which response is more helpful, harmless, and concise? Pick one.
GPT-4 Response
Response 2 is more helpful, harmless, and concise. It directly addresses the question about practicing the skill alone by suggesting a specific role-play scenario. It provides a clear example of how to disagree without causing an argument, which is precisely what the human asked for. It also avoids suggesting potentially more complex or less safe scenarios involving other people, as in Response 1, which could lead to real disagreements or conflicts if not handled carefully.
Appendix G Additional Experiments Results
G.1 Reward Distribution on HH Dataset
We also verify our method by comparing the reward distribution of responses. We train Qwen1.5-0.5B and Qwen1.5-1.8B models on the HH dataset using various methods. We then utilize trained models to generate responses according to the prompts within the test split of the HH dataset and score them with a reward model. As Figure 5 shows, all methods tend to increase the expected reward. However, our method achieves a greater increase, surpassing multiple methods including LiPO, DPOBT, DPOPL, and others.
Figure 5: Analysis of reward distribution on the test split of the HH dataset. We train Qwen1.5-0.5B (Left) and Qwen1.5-1.8B (Right) with various methods and compare the reward distribution of responses generated by trained models.
G.2 Additional DiffNDCG Discounts Experiments
To further analyze the impact of discount factors on our diffNDCG metric, we conducted experiments using various diffNDCG variants, including diffNDCG with Adaptive Rank Policy Score and with Policy Reference Ratio Score. We systematically evaluated the impact of varying discount factors across these diffNDCG variants and report results in Table 10. All experimental parameters and configurations followed the setup detailed in Section E.

### v2

@@ IV-C Differentiable Normalized Discounted Cumulative Gain Loss
$\mathrm{NDCG}(\hat{\mathbf{s}}_{\theta},\mathbf{s})=\frac{1}{\rm iDCG}\sum_{d=1}^{K}\frac{2^{\psi(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})}-1}{\log_{2}(1+d)}.$ (12)
The calculation of $\psi(d,\mathbf{s},\hat{\mathbf{s}}{\theta})$ involves sorting operations, making direct optimization of NDCG through gradient descent infeasible due to the non-differentiable nature of sorting. The sorting process can be formalized using a permutation matrix $\hat{\mathbf{s}}_{\theta}$, where each column $d$ indicates the position of each element of $\hat{\mathbf{s}}_{\theta}$ in the sorted order. Specifically, if the $j$-th element of $\hat{\mathbf{s}}_{\theta}$ is to be placed in the $d\rm{-th}$ position, then the entry $(j,d)$ in $\mathbf{P}_{\rm{hard}}$ is set to 1, and all other entries in the $d$-th column are set to 0. Thus, $\psi(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})$ can be expressed as $\psi(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})=[\mathbf{P}_{\rm{hard}}^{\top}\cdot\mathbf{s}]_{d}$. To address this issue, a differentiable permutation matrix $\mathbf{P}_{\rm soft}$ can be obtained through the differentiable sorting network based on $\hat{\mathbf{s}}_{\theta}$, as defined in Eq. 9. With $\mathbf{P}_{\rm soft}$, the relaxed, differentiable score $\psi^{\prime}(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})$ at ranking position $d$ can be computed as:
Finally, we consider the following optimization problem:
$\min_{\theta}\mathcal{L}_{\rm diff}=\frac{1}{|D|}\sum_{(x,\mathbf{y},\mathbf{s})\in D}\ell_{\rm diff}(\mathbf{M}_{\rm ad}(x,\mathbf{y},\pi_{\theta}),\mathbf{s}),$ (14)
where $\ell_{\rm diff}(\cdot,\cdot)=-\mathrm{diffNDCG}(\cdot,\cdot).$ The pseudo code of DRPO is presented in Algorithm 1.
TABLE I: Comparisons between our method and baselines. We report GPT-4 Win Rate (vs Chosen) and Reward Model Win Rate (vs Chosen and SFT). All Learning To Rank methods use PRR score.
Base Model Qwen1.5-0.5B Qwen1.5-1.8B
@@
Preference Alignment Methods
Learning To Rank Methods
Diff Sorting 28.97%($\pm$ 1.46) 44.53%($\pm$ 1.65) 66.41%($\pm$ 3.78) 49.39%($\pm$ 2.55) 64.06%($\pm$ 1.75) 78.56%($\pm$ 2.31)
V Experiments
V-A Experimental Settings
Datasets. Anthropic’s Helpful and Harmless (HH) [7] contains 161k/8.5k training/test samples. Each sample consists of a prompt and a pair of responses (chosen and reject), where “chosen" represents the preferred response and “reject" represents the less preferred response. We also generate additional responses for each prompt and rate each response using a reward model DeBERTa  [^2] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2, resulting in ranking preference data of a list size $K=8$. For more details, please refer to Appendix. UltraFeedback [14] contains 64k prompts. Each prompt corresponds to four responses, and every response has a score annotated by AI (e.g., GPT-4 and Gemini [63, 64]). VLFeedback [13] consists of 80k multi-modal samples from various sources.
Models. Our experiments are mainly based on Qwen1.5 model [49] with a range of parameters from 0.5B to 1.8B and Mistral model [48] with 7B parameter size. In addition, we also train Qwen-VL-Chat [65], a large-scale vision-language model, to evaluate the performance of our method on multi-modal preference alignment task [66].
Baseline Methods. To validate the effectiveness of our method, we conduct comparison experiments with representative baselines. In our experiments, we mainly compare our method with SFT, DPO [12], PRO [21], LiPO [15], DPOBT and DPOPL [12]. DPOBT adapts DPO to ranking preferences by decomposing ranked lists into pairwise comparisons, as proposed by [15]. Furthermore, DPOPL is proposed based on ranking preference data by [12], which leverages the Plackett-Luce preference model [32], a generalization of the Bradley-Terry model [67] that accommodates full rankings rather than just pairwise comparisons.
Furthermore, we implemented several differentiable sorting algorithms such as Fast Soft Sort [58], Neural Sort [47] and learning-to-rank methods, including ListNet [57], PiRank [46] for list preference alignment. Unless otherwise specified, the learning-to-rank methods calculate scores using the Policy Reference Ratio Score (PRR).
TABLE II: Performance comparison across different methods on Qwen models. RM and RM* represent DeBERTa and QRM-Llama3.1-8B-v2 reward models respectively. DRPO achieves superior performance across all evaluation metrics.
Method RM Score RM* Score MMLU GSM8K TruthfulQA
Qwen1.5-0.5B
SFT -1.7519 0.5204 39.25 14.78 38.64
PRO -1.4473 0.5449 36.19 14.85 38.26
Fast Soft Sort -0.7946 0.6109 38.82 18.45 37.77
LiPO -0.6350 0.6244 40.26 17.36 38.87
DRPO -0.5195 0.6263 41.34 18.65 39.51
Qwen1.5-1.8B
SFT -1.4651 0.5346 46.02 31.61 40.48
PRO -1.1162 0.5834 43.06 30.78 40.55
Fast Soft Sort -0.1291 0.6779 44.87 33.51 41.64
LiPO -0.0346 0.6891 46.57 33.88 42.41
DRPO 0.5372 0.7082 47.32 35.14 44.67
TABLE III: Performance comparison on LLama3-8B across different methods. We evaluate win rates (vs Chosen) and absolute scores using two reward models: RM (DeBERTa) and RM* (QRM-LLaMA3.1-8B-v2).
Method RM Win Rate RM Score RM* Win Rate RM* Score
LLama3-8B
SFT 44.33% -1.1930 53.12% 0.5921
LIRE 78.32% -0.0737 79.28% 0.6996
LiPO 82.14% 0.3142 83.72% 0.7121
DRPO 85.35% 0.5476 86.32% 0.7333
TABLE IV: Comparison of PRR replacement with ARP Score across ranking methods. While ARP improves the baseline methods, DRPO still maintains better performance.
@@ V-A Experimental Settings
Qwen1.5-0.5B
DPOBT(K=2) 28.26% 34.96%
DPOBT(K=2) + ARP 31.72%(+3.46) 43.55%(+8.59)
PiRank 26.08% 38.87%
PiRank + ARP 42.35%(+16.27) 57.03%(+18.16)
Fast Soft Sort 37.58% 51.95%
Fast Soft Sort + ARP 40.69%(+3.11) 54.49%(+2.54)
DRPO 42.80% 58.40%
Qwen1.5-1.8B
DPOBT(K=2) 47.60% 57.62%
DPOBT(K=2) + ARP 50.37%(+2.77) 61.37%(+3.75)
PiRank 50.71% 56.64%
PiRank + ARP 66.56%(+15.85) 75.69%(+18.95)
Fast Soft Sort 61.90% 71.87%
Fast Soft Sort + ARP 66.14%(+4.24) 77.92%(+6.05)
DRPO 69.08% 82.61%
Implementation details.
For experiments on the HH dataset, we use a learning rate of $5\times 10^{-7}$. For the UltraFeedback and VLFeedback datasets, we employ a learning rate of $5\times 10^{-6}$. Across all experiments, we adopt a global batch size of 32, a cosine learning rate schedule with a warmup ratio of 0.1, and the RMSprop optimizer [68]. All models are trained for one epoch on the entire training split. We set the maximum prompt length to 512 and the maximum context length to 1024. For LLaMA-3-8B [4] and Mistral-7B [^3] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta, we utilize QLoRA [69] to reduce memory consumption and facilitate efficient training. We configure our diffNDCG loss function with the following default hyperparameters: ranking constant $\tau=0.2$, coefficient $\beta=1$, and update rate $\theta=0.9999$ for the Adaptive Rank Policy Score (Eq.5), and steepness $\alpha=1$(Eq.8) for the differentiable sorting network. For DPO-style methods including DPO, DPOBT, and LiPO, we follow the same hyperparameter settings as in DPO [12], with $\beta=0.1$.
  • GPT-4 based Win Rate: we use GPT-4 to compare which of generated responses is more preferred, and measure performance by win rate of responses.
  • Open Benchmarks: We evaluate different methods on the UltraFeedback dataset using AlpacaEval2.0 [70] and MT-Bench [71] benchmarks. For the VLFeedback dataset, we employ MME [72], MMBench [73], and MMVet [74] benchmarks. Additionally, we assess performance on general benchmarks including MMLU [75], GSM8K [76], and TruthfulQA [77].
TABLE V: Ablation experiments of individual components.
@@
DRPO-w/o ada & diff 28.97%($\pm$1.46) 44.53%($\pm$1.65)
DRPO-w/o diff 33.06%($\pm$5.33) 47.66%($\pm$3.49)
DRPO-w/o ada 38.30%($\pm$3.35) 53.91%($\pm$1.46)
DRPO 42.80%($\pm$5.01) 58.40%($\pm$2.94)
V-B Experiments Results on Anthropic’s Helpful and Harmless Dataset
ă
Main Results. Experiments conducted on the HH dataset are presented in Table I, demonstrating the effectiveness of our method. Our method, DRPO, outperforms baseline methods including SFT, DPOBT, DPOPL, PRO, LiPO, and other LTR methods across different model scales. The GPT-4 Win Rate has an improvement of $5.22\%{\sim}6.13\%$, and Reward Model Win Rate has an improvement of $4.69\%{\sim}8.98\%$ and $0.78\%{\sim}2.73\%$. Several key observations emerge from these results. First, even without additional modifications, our Sorting Network outperforms most conventional sorting methods, such as Neural Sort. Second, directly extending existing alignment methods (e.g., DPO) to ranking list scenarios yields performance improvements. Third, while LiPO employs a metric weighting scheme to account for the relative importance of responses at different positions and achieves superior performance compared to DPOBT, DRPO leverages diffNDCG to precisely quantify response contributions at each ranking position, resulting in significantly enhanced performance over LiPO. Additionally, Appendix presents a comparison of reward distributions for model-generated responses, revealing improvements in response rewards when using our method.
TABLE VI: Effect of Discount Factor Selection on diffNDCG.
@@ V-B Experiments Results on Anthropic’s Helpful and Harmless Dataset
$1/{\sqrt{r}}$ 40.53%($\pm$4.36) 79.29%($\pm$2.89) 59.91%
$1/r$ 43.37%($\pm$5.14) 78.32%($\pm$2.94) 60.84%
$1/{r^{2}}$ 40.37%($\pm$4.17) 79.30%($\pm$3.38) 59.83%
$1/{\mathrm{log}(1+r)}$ 42.80%($\pm$5.01) 79.88%($\pm$3.72) 61.34%
TABLE VII: Time complexity and memory usage comparisons.
Method Time Complexity Run Time Memory
PRO $\mathcal{O}(L^{2})$ 0.2806s 24.57GB
LiPO $\mathcal{O}(L^{2})$ 0.3269s 24.08GB
Neural Sort $\mathcal{O}(L^{2})$ 0.2585s 24.01GB
Pirank $\mathcal{O}(L^{2})$ 0.2635s 25.80GB
Fast Soft Sort $\mathcal{O}(L\log(L))$ 0.2618s 23.21GB
ListNet $\mathcal{O}(L^{2})$ 0.2652s 21.10GB
DRPO (odd-even) $\mathcal{O}(L^{2})$ 0.2641s 23.39GB
DRPO (bitonic) $\mathcal{O}(L\log^{2}(L))$ 0.2560s 23.43GB
Validation on Multiple Reward Models and Open Benchmarks. To comprehensively validate our method, we employ multiple reward models and evaluation metrics. Specifically, we use QRM-Llama3.1-8B-v2 [78], a top-performing model in RewardBench [79], alongside the DeBERTa-based reward model. We evaluate on three established benchmarks: MMLU [75], GSM8K [76], and TruthfulQA [77].
Table II presents comprehensive comparisons with competitive baselines. "RM" and "RM*" denote evaluations using DeBERTa and QRM-Llama3.1-8B-v2 respectively, while "Score" represents the absolute reward score. DRPO consistently outperforms all baselines across all metrics, with consistent preferences demonstrated by both reward models, validating the effectiveness of our approach.
Comparison on larger LLMs. We validate our method on a larger-scale LLM (LLaMA3-8B [4]). We further compare our approach with an additional ranking-based method LIRE [80] and a strong baseline from previous experiments, LiPO [15]. Results are presented in Table III. While LIRE and LiPO both improve over the SFT baseline, our method consistently outperforms all competing approaches across different model scales. This demonstrates the effectiveness and scalability of our approach regardless of model size.
V-C Ablation Studies
Ablation Studies on Adaptive Rank Policy Score. To assess the efficacy of our proposed Adaptive Rank Policy Score, we conducted a comparative analysis by integrating it into multiple baseline models. The results are presented in Table IV. Our experimental results demonstrate that the Adaptive Rank Policy Score significantly enhances baseline model performance across various sizes, showcasing the advantages of applying ARP over the PRR score.
Furthermore, we conduct an ablation study on our method by progressively removing Adaptive Rank Policy Score(ARP) and diffNDCG, denoted as ’w/o ARP’ and ’w/o diffNDCG’ respectively. As shown in Table V, we observe that directly using the cross-entropy loss between the ground truth permutation matrix $\mathbf{P}_{\rm ground}$ and the predicted permutation matrix $\mathbf{P}_{\rm soft}$, without ARP and diffNDCG, leads to performance degradation. This fails to account for the varying importance of responses at different ranking positions. Additionally, when the ARP is replaced by PRR in Eq (2), performance decreases by $4.09\%-4.5\%$ and 3.13%-4.49% for GPT-4 Win Rate and RM Win Rate, respectively.
Ablation Studies on DiffNDCG Discounts Factors.
To further analyze the impact of discount factors on diffNDCG, we conducted ablation studies by comparing our default discount factors with alternative variants. Specifically, we set up three variants for comparison: inverse square root ($1/\sqrt{r}$), inverse linear ($1/r$), and inverse quadratic ($1/r^{2}$), where $r$ denotes the ranking position. Our default formulation employs the logarithmic discount factor $1/\log(1+r)$. The experimental results are presented in Table VI.